# Intermediate Relation Size Bounds for Select-Project-Join Query Plans: Asymptotically Tight Characterizations

### v1

@@ 4. Relating queries to structures
[…]
Proof.
We show, by induction on the structure of $p$, how to derive an extended $p$-representation $(\mathbf{A},\overline{a},\approx)$ and $p$-decomposition thereof which satisfies the containment property. We then have that $(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle)$ is a $p$-representation along with a $p$-decomposition. We prove the statement inductively over the structure of $p$. Here, we show the case where $p$’s outer-most operator is a join; the other cases are relatively straightforward and treated in the appendix.
Suppose that $p$ has the form $\Join_{\theta}(p_{1},\ldots,p_{\ell})$, where each $p_{i}$ has arity $m_{i}$, for $i\in[\ell]$, and $p$ has arity $m=m_{1}+\cdots+m_{\ell}$. By induction, there exists (for each $i$) an extended $p_{i}$-representation $(\mathbf{A}_{i},\overline{t_{i}},\approx_{i})$. We may assume without loss of generality that the universes $\mathsf{dom}(\mathbf{A}_{i})$ are pairwise disjoint (if they are not, their elements can be renamed to achieve this). We let $\overline{a}=(a_{1},\ldots,a_{m})$ denote the tuple $\overline{t_{1}}\ldots\overline{t_{\ell}}$, that is, the tuple obtained by concatenating the entries of the tuples $\overline{t_{i}}$, $i\in[\ell]$, in order. Define $\approx_{0}$ as the relation $\bigcup_{i=1}^{\ell}\approx_{i}$, i.e., collecting all previously known equivalences, and define $\approx$ as the relation $\approx_{0}\cup\bigcup_{(j=k)\in\theta}\{(a_{j},a_{k})\}$, i.e., adding the equivalences newly introduced by $\theta$. We define $\mathbf{A}$ as $\mathbf{A}_{1}\cup\cdots\cup\mathbf{A}_{\ell}$. Then we have that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation, since for any structure $\mathbf{D}$ over $\sigma$:
$\mathsf{homs}(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle,\mathbf{D})$
$=\sigma_{\theta}(\mathsf{homs}(\mathbf{A}\langle\approx_{0}\rangle,\overline{a}\langle\approx_{0}\rangle,\mathbf{D}))$
$=\sigma_{\theta}(\mathsf{homs}(\mathbf{A}_{1}\langle\approx_{1}\rangle,\overline{t_{1}}\langle\approx_{1}\rangle,\mathbf{D})\times\cdots\times\mathsf{homs}(\mathbf{A}_{\ell}\langle\approx_{\ell}\rangle,\overline{t_{\ell}}\langle\approx_{\ell}\rangle,\mathbf{D}))$
$=\sigma_{\theta}(\mathsf{out}(p_{1},\mathbf{D})\times\cdots\times\mathsf{out}(p_{\ell},\mathbf{D}))$
$=\mathsf{out}(p,\mathbf{D}).$
Here, the first equality comes from our definition of $\approx$ from $\approx_{0}$, and the fact that each pair in $\approx_{0}$ has elements contained entirely in a universe $\mathsf{dom}(\mathbf{A}_{i})$; the second equality follows from this fact and our definitions of $\overline{a}$ and $\approx_{0}$; the third equality follows from our assumption that each $(\mathbf{A}_{i},\overline{t_{i}},\approx_{i})$ is an extended $p_{i}$-representation; the last equality follows from the definition of $\mathsf{out}$.
By induction, each given extended $p_{i}$-representation has a $p_{i}$-decomposition $E_{i}=(T_{i},\chi_{i})$, where $r_{i}$ is the root node of $T_{i}$, with associated mappings $\alpha_{i},\beta_{i}$. We assume without loss of generality that the vertex sets $V(T_{i})$ are pairwise disjoint (if not, they can be renamed to achieve this). We provide a $p$-decomposition as follows. Combining the trees $T_{i}$, create a rooted tree $T$ with one extra vertex $r$ whose children are $r_{1},\ldots,r_{\ell}$. Define $\alpha:\mathsf{subplans}(p)\to V(T)$ to be the mapping that extends each mapping $\alpha_{i}$ and where $\alpha(p)=r$; define $\beta:\mathsf{subplans}(p)\to\mathsf{dom}(\mathbf{A})\langle\approx\rangle^{*}$ to be the mapping with $\beta(p)=\overline{a}\langle\approx\rangle$, and where, for each subplan $q$ of a plan $p_{i}$, when $\beta_{i}(q)=(c_{1},\ldots,c_{m_{i}})$, $\beta(q)$ is defined as the tuple obtained by further dividing each entry by $\approx^{*}$, that is, as the tuple $(c^{\prime}_{1},\ldots,c^{\prime}_{m_{i}})$ where each $c^{\prime}_{j}$ is the unique $\approx^{*}$-equivalence class containing $c_{j}$; such a class exists, since each $c_{j}$ is a $\approx_{i}^{*}$-equivalence class, and $\approx_{i}^{*}$ is contained in $\approx^{*}$ (due to $\approx_{i}$ being contained in $\approx$).
Define $\chi(r)=\{\overline{a}\langle\approx\rangle\}$. For each subplan $q$ of $p$ with $q\neq p$, define $\chi({\alpha(q)})=\{\beta(q)\}$. It is straightforward to verify that $(T,\chi)$ is a rooted tree decomposition as required by the definition of $p$-decomposition with mappings $\alpha$ and $\beta$ as defined above.
It remains to show that the $p$-decomposition $(T,\chi)$ satisfies the containment property. We showed that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation, so it holds in the case that $q=p$. Suppose that $q$ is a subplan of $p$ with $q\neq p$. Then $q$ is a subplan of a plan $p_{j}$, for some $j\in[\ell]$. Since $(\mathbf{A}_{j},\overline{t_{j}},\approx_{j})$ is an extended $p_{j}$-representation, we have for any structure $\mathbf{D}$ over $\sigma$ that
$\mathsf{out}(q,\mathbf{D})\supseteq\mathsf{homs}(\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}},\beta_{j}(q),\mathbf{D}).$
Let us make some observations. First, we have $\alpha(q)=\alpha_{i}(q)$, and so $\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}}$ is a substructure of $\mathbf{A}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}}$. We can write $\beta_{j}(q)$ as a tuple $(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx_{j}\rangle$ where $a^{\prime}_{1},\ldots,a^{\prime}_{k}\in\mathsf{dom}(\mathbf{A}_{j})$. By our definition of $\beta(q)$ from $\beta_{j}(q)$, we have $\beta(q)=(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle$. We have
$\mathsf{homs}(\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}},\beta_{j}(q),\mathbf{D})$
$=\mathsf{homs}(\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}},(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx_{j}\rangle,\mathbf{D})$
$\supseteq\mathsf{homs}(\mathbf{A}\langle\approx\rangle_{\leq\alpha(q),E},(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle,\mathbf{D}).$
The equality follows from the above observations on $\beta_{j}(q)$. The containment can be explained as follows. We have that $\approx_{j}$ is a subset of $\approx$, and, on nodes below $\alpha(q)$, the tree decomposition $E$ is obtained from the tree decomposition $E_{j}$ by further dividing each element in a set associated via $\chi_{j}$ by $\approx^{*}$. We thus have a canonical homomorphism from $\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}}$ to $\mathbf{A}\langle\approx\rangle_{\leq\alpha(q),E}$, namely, the mapping that further divides each element of the first structure by $\approx^{*}$. Since this homomorphism maps the tuple $(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx_{j}\rangle$ to the tuple $(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle$, the containment holds by Proposition 2.2.
By chaining together the established relationships and using $\beta(q)=(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle$, we conclude $\mathsf{out}(q,\mathbf{D})\supseteq\mathsf{homs}(\mathbf{A}\langle\approx\rangle_{\leq\alpha(q),E},\beta(q),\mathbf{D})$, as required.
We finish this section by showing how the $p$-representation of a plan $p$ can be used to determine the output degree of the plan. To this end, we present two lemmas providing upper and lower bounds on the size of the result of evaluating an open structure on a structure. This lemma, providing an upper bound, is a reformulation of Theorem 4.4 of (Gottlob et al., 2012).
Lemma 4.5.
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $\mathbf{D}$ be a structure over $\sigma$ that satisfies $\Sigma$. Let $(\mathbf{A},\overline{a})$ be an open structure over $\sigma$. Let $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$. It holds that
$|\mathsf{homs}(\mathbf{A},\{\overline{a}\},\mathbf{D})|\leq M_{\mathbf{D}}^{C_{\Sigma}^{\mathbf{B}}(\{\overline{b}\})}.$
We next present a complementary lemma that provides the matching lower bound and which again follows directly from Theorem 4.4 of (Gottlob et al., 2012).
Lemma 4.6.
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $\mathbf{A}$ be a structure over $\sigma$. For any open structure $(\mathbf{A},\overline{a})$, there exists a constant $c>0$ such that, with $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$, there exists a sequence $(\mathbf{D}_{n})_{n\geq 1}$ of structures over $\sigma$ that satisfy $\Sigma$, with $M_{\mathbf{D}_{n}}$ unbounded, and where, for each $n~{}\geq~{}1$,
$|\mathsf{homs}(\mathbf{A},\{\overline{a}\},\mathbf{D}_{n})|\geq cM_{\mathbf{D}_{n}}^{C_{\Sigma}^{\mathbf{B}}(\{\overline{b}\})}.$
Now, Theorem 4.4 guarantees the existence of a $p$-representation for any SPJ plan $p$. Combining the existence of a $p$-representation with the upper and lower bounds given by Lemmas 4.5 and 4.6, respectively, we yield the following proposition pin-pointing the output degree of a plan $p$.
Proposition 4.7.
5. Intermediate Size Bounds
@@
We devote this section to formalizing and establishing the statements used to prove Theorem 3.3. We start with statement (i), i.e. we show that we can synthesize a plan which is $\Sigma$-equivalent to the one given originally and which has the desired intermediate degree.
@@
There exists a polynomial-time algorithm that, given a set $\Sigma$ of unary keys over a signature $\sigma$ and an open structure $(\mathbf{C},\overline{c})$ over $\sigma$ with $\textrm{Chase}_{\Sigma}(\mathbf{C},\overline{c})=(\mathbf{C},\overline{c})$ and which is a $p$-representation for some plan $p$, outputs a well-behaved plan $p^{\prime}$ where (a) $p^{\prime}$ is $\Sigma$-semantically equivalent to $p$, and (b) $p^{\prime}$ has intermediate degree $\leq C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$.
Proof.
The heart of the proof can be derived from argumentation in Section 4 of (Gottlob et al., 2012). As we draw on this section, we employ the terminology of (Gottlob et al., 2012), and refer to their argumentation. In particular, their result refers to conjunctive queries and functional dependencies. The former are known to be equivalent to SPJ queries (and are in syntax close to our $p$-representations), while the latter generalize key constraints. In the following arguments, as in (Gottlob et al., 2012), for elements $X_{i},X_{j}\in\mathsf{dom}(\mathbf{C})$, we use the notation $X_{i}\rightarrow X_{j}$ to say that element $X_{j}$ is functionally determined by $X_{i}$ via some functional dependency. In Section 4 of (Gottlob et al., 2012), it is explained how to transform a conjunctive query $Q_{c}$, to which the chase has been applied, to a conjunctive query $Q^{\prime}$ over a modified set of relations that has no functional dependencies. Let us for the moment put aside the output relation and output variables, and consider the action of this transformation on the input relations and variables. The transformation considers functional dependencies of the form $X_{i}\rightarrow X_{j}$ over pairs of input variables, and iteratively removes them; a removal is performed by adding entailed functional dependencies, and by expanding each atom including $X_{i}$ to a new atom where $X_{j}$ is added (if not already present).
For each functional dependency $X_{i}\rightarrow X_{j}$, we can maintain a binary relation $S_{ij}$ whose first coordinate is a key for the relation, and such that adding $S_{ij}(X_{i},X_{j})$ to the query atoms would not change the query answers; this holds initially by taking $S_{ij}$ to be an appropriate projection of the relation of an atom mentioning both $X_{i}$ and $X_{j}$, and then when an entailed dependency $X_{k}\rightarrow X_{j}$ is added, due to dependencies $X_{k}\rightarrow X_{i}$ and $X_{i}\rightarrow X_{j}$, we can derive $S_{kj}$ (if needed) by taking the natural join of $S_{ki}$ and $S_{ij}$ and then projecting the middle coordinate. It is readily verified that the sizes of these relations $S_{ij}$ never exceeds the size of any input relation.
When an atom including a variable $X_{i}$ is expanded to include the variable $X_{j}$, a new relation for the atom can be computed by joining the original relation with the relation $S_{ij}$. Since each relation $S_{ij}$ expresses a functional dependency, this never increases the size of any relation.
The result of the transformation, applied to the input atoms of a conjunctive query, is thus a new set of input atoms where every new relation can be derived from the original relations via a well-behaved SPJ plan of intermediate degree $1$, where every new relation has size bounded above by an original relation, and where the answers (over all variables) to the new set of atoms is the same as the answers to the original set. Consider the application of the transformation to the input atoms of the conjunctive query corresponding to $\mathbf{C}$. Let $\mathbf{C}^{\prime}$ be the relational structure representing the new set of atoms. It is moreover shown in Lemma 4.7 of (Gottlob et al., 2012) that the color number of an overall query (including input variables) is not changed by this transformation, and this implies that $C^{\mathbf{C}^{\prime}}_{\Sigma}(S)\leq C^{\mathbf{C}^{\prime}}_{\Sigma}(S^{\prime})=C^{\mathbf{C}}_{\Sigma}(S)$ for any set $S$ of variables and its expansion $S^{\prime}$ after applying the transformation. Let us then consider $(\mathbf{C}^{\prime},\overline{c})$; due to the just-given inequality, we have $C_{\Sigma}^{\mathbf{C}^{\prime}}\textup{-width}(\mathbf{C}^{\prime},\overline{c})\leq C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$, and it thus suffices to establish the result for the empty set of unary keys $\Sigma^{\prime}=\emptyset$ along with $(\mathbf{C}^{\prime},\overline{c})$, in place of $\Sigma$ with $(\mathbf{C},\overline{c})$.
To establish the result for $(\mathbf{C}^{\prime},\overline{c})$ in the absence of keys, it suffices to compute a tree decomposition $F$ of $(\mathbf{C}^{\prime},\overline{c})$ where $C_{\Sigma}^{\mathbf{C}^{\prime}}\textup{-width}(F)$ is equal to $C_{\Sigma}^{\mathbf{C}^{\prime}}\textup{-width}(\mathbf{C}^{\prime},\overline{c})$, which can be done in exponential time, and then compute a plan that materializes the answers to each bag, where the answers for a bag with variables $S$ would be considered to be the join of the set of atoms obtained from $\mathbf{C}^{\prime}$’s conjunctive query by taking each atom, and projecting onto $S$. These different sets of answers can then be combined via joins and projects in a standard way, following the structure of the tree decomposition. But to find a plan that does this materialization for a bag with variables $\{a_{1},\ldots,a_{n}\}$, we simply follow the proof of Theorem 6 of (Atserias et al., 2013), which explains how to construct a join-project plan for any pure join plan; in doing so, they define plans $\phi_{1},\ldots,\phi_{n}$; each of these plans is implementable as a well-behaved one, since they iteratively introduce one variable at a time: in their notation, the plan $\phi_{i}$, which can be viewed as a multiway join, concerns variables $\{a_{1},\ldots,a_{i}\}$.
We continue to show that statement (ii) of Section 3 also holds. To that end, let us first formalize the statement. In essence, it provides a lower bound on the intermediate degree of any plan that is $\Sigma$-semantically equivalent to $p$.
@@
Proof.
We show how to pass from any map in $\mathsf{homs}(\mathbf{C},S~{}\cap~{}\mathsf{dom}(\mathbf{C}),\mathbf{D})$ to a map in $\mathsf{homs}(\mathbf{B},S,\mathbf{D})$, and then explain why the passage is injective, which suffices. Suppose that $g\in\mathsf{homs}(\mathbf{C},S\cap\mathsf{dom}(\mathbf{C}),\mathbf{D})$ and $h$ is a retraction from $\mathbf{B}$ to $\mathbf{C}$. Then there exists an extension $g^{\prime}$ of $g$ that is in $\mathsf{homs}(\mathbf{C},\mathsf{dom}(\mathbf{C}),\mathbf{D})$, from which it follows that the composition $g^{\prime}\circ h$ is in $\mathsf{homs}(\mathbf{B},\mathsf{dom}(\mathbf{B}),\mathbf{D})$ (in writing this composition, $h$ is applied first). It follows that the restriction $(g^{\prime}\circ h)\upharpoonright S$ is in $\mathsf{homs}(\mathbf{B},S,\mathbf{D})$. This passage is injective since if two mappings in $\mathsf{homs}(\mathbf{C},S\cap\mathsf{dom}(\mathbf{C}),\mathbf{D})$ differ on an element $c$ of $S\cap\mathsf{dom}(\mathbf{C})$, the two mappings that they are passed to will also differ on $c$, due to the assumption that $h$ acts as the identity on $C$.
We proceed to prove our final theorem.
Proof of Theorem 5.2.
Fix a set $\Sigma$ of unary keys and a SPJ plan $p$ over $\Sigma$. Let $(\mathbf{A},\overline{a})$ be a $p$-representation, $(\mathbf{B},\overline{b})=\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, and $(\mathbf{C},\overline{c})$ be a core of $(\mathbf{B},\overline{b})$. For any structure $\mathbf{D}$ satisfying $\Sigma$ we have that
$\mathsf{out}(p,\mathbf{D})=\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})=\mathsf{homs}(\mathbf{B},\overline{b},\mathbf{D})=\mathsf{homs}(\mathbf{C},\overline{c},\mathbf{D}),$
where the first equality holds by definition of $p$-representation, the second equality holds by definition of the chase procedure and the third equality holds by Propositions 2.4.
Let $p^{+}$ be any SPJ plan that is $\Sigma$-semantically equivalent to $p$. Let $(\mathbf{A}^{+},\overline{a^{+}})$ be a $p^{+}$-representation and $(\mathbf{B}^{+},\overline{b^{+}})=\textrm{Chase}_{\Sigma}(\mathbf{A}^{+},\overline{a^{+}})$. For any structure $\mathbf{D}$ satisfying $\Sigma$ we have that
$\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p^{+},\mathbf{D})=\mathsf{homs}(\mathbf{A}^{+},\overline{a^{+}},\mathbf{D})=\mathsf{homs}(\mathbf{B}^{+},\overline{b^{+}},\mathbf{D}),$
where the first equality holds by $\Sigma$-semantic equivalence and the second and third equality hold by definition as above.
Since we have that $\mathsf{homs}(\mathbf{C},\overline{c},\mathbf{D})=\mathsf{homs}(\mathbf{B}^{+},\overline{b^{+}},\mathbf{D})$ for all structures $\mathbf{D}$ satisfying $\Sigma$, and we know that both $\mathbf{C}$ and $\mathbf{B}^{+}$ are structures satisfying $\Sigma$, it holds specifically that $\overline{c}\in\mathsf{homs}(\mathbf{C},\overline{c},\mathbf{C})=\mathsf{homs}(\mathbf{B}^{+},\overline{b^{+}},\mathbf{C})$ and $\overline{b^{+}}\in\mathsf{homs}(\mathbf{B}^{+},\overline{b^{+}},\mathbf{B}^{+})=\mathsf{homs}(\mathbf{C},\overline{c},\mathbf{B}^{+})$. This immediately implies that there is a homomorphism from $(\mathbf{B}^{+},\overline{b^{+}})$ to $(\mathbf{C},\overline{c})$, and vice-versa, showing that the two open structures are homomorphically equivalent. Since $(\mathbf{C},\overline{c})$ is a core, by Proposition 2.4 we have that $\mathsf{aug}(\mathbf{C},\overline{c})$ is isomorphic to a substructure of $\mathsf{aug}(\mathbf{B}^{+},\overline{b^{+}})$. By renaming the elements of $\mathsf{dom}(\mathbf{B}^{+})$, we can achieve that $\mathsf{aug}(\mathbf{C},\overline{c})$ is a core of (and thus a substructure of) $\mathsf{aug}(\mathbf{B}^{+},\overline{b^{+}})$. We then have that $\overline{b^{+}}=\overline{c}$.
By Theorem 4.4, there exists a $p^{+}$-decomposition $E^{+}=(T,\chi^{+})$ of $(\mathbf{A}^{+},\overline{a^{+}})$ with associated mappings $\alpha,\beta$. Within $E^{+}$, we equate elements following the chase procedure on $(\mathbf{A},\overline{a})$ such that the resulting $E^{+}$ is a decomposition of $(\mathbf{B}^{+},\overline{b^{+}})$. We define a tree decomposition of $(\mathbf{C},\overline{c})$ as follows. Take the pair $(T,\chi)$ where $T$ is as in $E^{+}$, and where, for each vertex $t$ of $T$, we define $\chi(t)=\chi^{+}(t)\cap\mathsf{dom}(\mathbf{C})$. It is straightforward to verify that $(T,\chi)$ is a tree decomposition of $(\mathbf{C},\overline{c})$. By Lemma 5.3, there exists a vertex $t$ of $T$, a constant $c^{\prime}>0$, and structures $(\mathbf{D}_{n})_{n\geq 1}$ with $M_{\mathbf{D}_{n}}$ unbounded, where $|\mathsf{homs}(\mathbf{C},\chi(t),\mathbf{D}_{n})|\geq c^{\prime}M^{d}_{\mathbf{D}_{n}}$. There exists a subplan $q^{+}$ of $p^{+}$ such that $\alpha(q^{+})=t$. We then have
$|\mathsf{out}(q^{+},\mathbf{D}_{n})|$ $\geq|\mathsf{homs}(\mathbf{B}_{\leq t,E^{+}},\chi^{+}(t),\mathbf{D}_{n})|$
$\geq|\mathsf{homs}(\mathbf{B},\chi^{+}(t),\mathbf{D}_{n})|$
$\geq|\mathsf{homs}(\mathbf{C},\chi(t),\mathbf{D}_{n})|$
$\geq c^{\prime}M^{d}_{\mathbf{D}_{n}}.$
The first inequality holds by the definition of $p^{+}$-decomposition; the second inequality holds since $\mathbf{B}_{\leq t,E^{+}}$ is a substructure of $\mathbf{B}$; the third inequality follows from Lemma 5.4, whose assumptions hold because $\mathsf{aug}(\mathbf{C},\overline{c})$ is a core of $\mathsf{aug}(\mathbf{B},\overline{b})$; the last inequality follows from our choice of $(\mathbf{D}_{n})_{n\geq 1}$. It follows that the subplan $q^{+}$ has output degree $\geq d$.
6. Conclusion

### v2

@@ 4.1 Relating Queries to Structures
4.2 Intermediate Size Bounds
@@
We devote this subsection to formalizing and establishing the statements used to prove Theorem 3.4. We start with statement (i), i.e. we show that we can synthesize a plan which is $\Sigma$-equivalent to the one given originally and which has the desired intermediate degree.
@@
There exists an exponential-time algorithm that, given a set $\Sigma$ of unary keys over a signature $\sigma$ and an open structure $(\mathbf{C},\overline{c})$ over $\sigma$ with $\textrm{Chase}_{\Sigma}(\mathbf{C},\overline{c})=(\mathbf{C},\overline{c})$ and which is a $p$-representation for some plan $p$, outputs a well-behaved plan $p^{\prime}$ where (a) $p^{\prime}$ is $\Sigma$-semantically equivalent to $p$, and (b) $p^{\prime}$ has intermediate degree $\leq C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$.
To show the above theorem, we combine proof techniques from [2] and [10]. As a first step, following arguments from Section 4 of [10], we transform the open structure $(\mathbf{C},\overline{c})$ together with $\Sigma$ to an open structure $(\mathbf{C}^{\prime},\overline{c})$ (over a modified signature $\sigma^{\prime}$) which has no constraints posed over it. This transformation can be achieved via the application of a well-behaved SPJ plan of intermediate degree 1, and where every new relation is of a size that is bounded from above by an original relation. Crucially, for any structure $\mathbf{D}$ satisfying $\Sigma$, we have that $(\mathbf{C},\overline{c})$ returns the same answers on $\mathbf{D}$ as $(\mathbf{C}^{\prime},\overline{c})$ on an analogous transformation of $\mathbf{D}$. Moreover, we have that $C_{\emptyset}^{\mathbf{C}^{\prime}}\textup{-width}(\mathbf{C}^{\prime},\overline{c})\leq C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$ and it thus suffices to show the result for the empty set of unary keys $\Sigma^{\prime}=\emptyset$ along with $(\mathbf{C}^{\prime},\overline{c})$, in place of $\Sigma$ with $(\mathbf{C},\overline{c})$. Finally, we compute a tree decomposition $F$ of $(\mathbf{C}^{\prime},\overline{c})$ where $C_{\Sigma^{\prime}}^{\mathbf{C}^{\prime}}\textup{-width}(F)$ is equal to $C_{\Sigma^{\prime}}^{\mathbf{C}^{\prime}}\textup{-width}(\mathbf{C}^{\prime},\overline{c})$, which can be done in exponential time, and then compute a SPJ plan $p^{\prime}$ following the tree composition using standard techniques from [2]. We provide a detailed proof in the appendix. We continue to show that statement (ii) also holds. To that end, let us first formalize the statement. In essence, it provides a lower bound on the intermediate degree of any plan that is $\Sigma$-semantically equivalent to $p$.
@@
We proceed to prove the final statement needed to show Theorem 3.4.
Proof 4.11 (Proof of Theorem 4.8).
[…]
5 SPJU Plans
Building on the established understanding of intermediate degree for SPJ plans from the previous section, we are now ready to prove Theorem 3.3. It is well-known that each SPJU plan $p$ is $\Sigma$-semantically equivalent to some plan $p^{\prime}$ which is a union $p_{1}\cup\cdots\cup p_{m}$ of SPJ plans. Let us note that the output degree of an SPJU plan can then be characterized as follows: for such a $\Sigma$-semantically equivalent plan $p^{\prime}$, the output degree of $p$ is equal to the maximum output degree over the plans $p_{1},\ldots,p_{m}$. To argue about the intermediate degree of an SPJU plan, we first give a theorem showing that a natural such translation to a union of SPJ plans exists and can be computed in exponential time, which has the additional property that the relation computed by a subplan of one of the SPJ plans is a subset of some relation computed by a subplan of the original SPJU plan; this inclusion then facilitates proving lower bounds on the intermediate degree of the original SPJU plan.
Theorem 5.1.
There exists an exponential-time algorithm that, given an SPJU plan $p$, outputs a sequence of SPJ plans $p_{1},\ldots,p_{m}$, a sequence of maps $\alpha_{1},\ldots,\alpha_{m}$, and a sequence of open structures $(\mathbf{A}_{1},\overline{a_{1}}),\ldots,(\mathbf{A}_{m},\overline{a_{m}})$ such that the following holds: 1. $p$ is semantically equivalent to $p_{1}\cup\cdots\cup p_{m}$, 2. for each $i=1,\ldots,m$, we have that $\alpha_{i}:\mathsf{subplans}(p_{i})\to\mathsf{subplans}(p)$ is a map such that (for each structure $\mathbf{D}$) when $q\in\mathsf{subplans}(p_{i})$, the containment $\mathsf{out}(q,\mathbf{D})\subseteq\mathsf{out}(\alpha_{i}(q),\mathbf{D})$ holds, 3. for each $i=1,\ldots,m$, the open structure $(\mathbf{A}_{i},\overline{a_{i}})$ is a $p_{i}$-representation, and $i\neq j$ implies that there is no homomorphism from $(\mathbf{A}_{i},\overline{a_{i}})$ to $(\mathbf{A}_{j},\overline{a_{j}})$.
On a high level, Theorem 3.3 can then be proved as follows. For some SPJU plan $p$, the algorithm first computes this convenient set of SPJ plans $p_{1},\ldots,p_{m}$ via Theorem 5.1. Then, for each obtained plan $p_{i}$, it computes a core $(\mathbf{C}_{i},\overline{c_{i}})$ of the chase of a $p$-representation, and then applies Theorem 3.4 to obtain a value $d_{i}$, and a plan $p^{\prime}_{i}$ such that $p_{i}$, $d_{i}$, and $p^{\prime}_{i}$ satisfy the properties given in that theorem statement. The algorithm outputs the value $d=\max_{i\in[m]}d_{i}$ and the plan $p^{\prime}=p^{\prime}_{1}\cup\cdots\cup p^{\prime}_{m}$. Since this algorithm outputs a union of well-behaved SPJ plans, also the resulting SPJU plan is well-behaved. It is easy to see that $p^{\prime}$ has intermediate degree $d$, it just remains to show that $d$ is indeed the SPJU-best possible intermediate degree. In essence, we prove that for any SPJU $q^{\prime}$ that is semantically equivalent to $p^{\prime}$, we have that from the sequence of SPJ plans $q^{\prime}_{1},\ldots,q^{\prime}_{n}$ obtained by applying Theorem 5.1 to $q^{\prime}$, there exists an index $\ell\in[n]$ such that $q^{\prime}_{\ell}$ does not have (for any $\epsilon>0$) intermediate degree $\leq(d-\epsilon)$ implying by condition (2) of Theorem 5.1 that $q^{\prime}$ does not have (for any $\epsilon>0$) intermediate degree $\leq(d-\epsilon)$. Full details can be found in the appendix.
6 Conclusion